OpenStudy (babyslapmafro):

Please help me simplify the following... (sqrt2)+(sqrt8)

OpenStudy (babyslapmafro):

\[\sqrt{2}+\sqrt{8}\]

OpenStudy (ksaimouli):

what is sqrt of 8

OpenStudy (anonymous):

\[\sqrt{8}=\sqrt{(2*2)*2}=2*\sqrt{2}\]

OpenStudy (ksaimouli):